# Random-Set Neural Networks (RS-NN)

### v2

@@ Abstract
Machine learning is increasingly deployed in safety-critical domains where robustness against adversarial attacks is crucial and erroneous predictions could lead to potentially catastrophic consequences. This highlights the need for learning systems to be equipped with the means to determine a model’s confidence in its prediction and the epistemic uncertainty associated with it, ‘to know when a model does not know’. In this paper, we propose a novel Random-Set Neural Network (RS-NN) for classification. RS-NN predicts belief functions rather than probability vectors over a set of classes using the mathematics of random sets, i.e., distributions over the power set of the sample space. RS-NN encodes the ‘epistemic’ uncertainty induced in machine learning by limited training sets via the size of the credal sets associated with the predicted belief functions. Our approach outperforms state-of-the-art Bayesian (LB-BNN, BNN-R) and Ensemble (ENN) methods in a classical evaluation setting in terms of performance, uncertainty estimation and out-of-distribution (OoD) detection on several benchmarks (CIFAR-10 vs SVHN/Intel-Image, MNIST vs FMNIST/KMNIST, ImageNet vs ImageNet-O) and scales effectively to large-scale architectures such as WideResNet-28-10, VGG16, Inception V3, EfficientNetB2, and ViT-Base.
@@ 1 Introduction
Despite its success, artificial intelligence, in its current form, struggles to provide reliable predictions when confronted with unfamiliar data guo2017calibration . This includes noisy samples papernot2016limitations deliberately designed to deceive models, or out-of-distribution data (OoD) beyond the model’s training distribution. This emphasizes the need for learning systems to be able to determine a model’s confidence in its own predictions, acknowledge the limits of its knowledge and gauge these limitations to make informed decisions by modelling the epistemic uncertainty associated with it, in essence, ‘to know when a model does not know’. Epistemic uncertainty pertains to uncertainty associated with our ignorance about the underlying data generation process kendall2017uncertainties ; hullermeier2021aleatoric ; manchingal2022epistemic . Within machine learning, a major source of epistemic uncertainty arises from the limited representativeness of the available training data, constrained in both quantity and quality.
Figure 1: Inference in a Bayesian Neural Network (top) as opposed to a Random-Set Neural Network (bottom), with corresponding measures of uncertainty and their sources.
Figure 2: Confidence scores of RS-NN and CNN for FGSM attack ($\epsilon$ = 0.05) on CIFAR-10.
@@
In this paper, we propose to model epistemic uncertainty using a random set approach Molchanov_2017 . As we focus on classification involving a finite list of classes, we model uncertainty using belief functions Shafer76 , the finite incarnation of random sets cuzzolin2018visions . The theory of belief functions, or Dempster-Shafer theory of evidence Shafer76 , is, in fact, a generalisation of Bayesian inference smets86bayes . Classical (discrete) probabilities are simply special belief functions, and Bayes’ rule is a special case of Dempster’s rule originally proposed for the combination of belief functions. Fig. 2 contrasts the inference processes of a Random-Set Neural Network (RS-NN) and a Bayesian Neural Network DBLP:journals/corr/abs-2007-06823 . In hierarchical Bayesian inference, the predictive posterior distribution is a second-order probability distribution hullermeier2021aleatoric . It involves sampling from predicted first-order distributions for a collection of possible models within the Bayesian Neural Network. A single (mean) prediction is derived by Bayesian Model Averaging (BMA), while uncertainty is measured by the entropy of the mean prediction and the variance of the posterior distributions.
@@
In a Random-Set NN (RS-NN) (Fig. 2, bottom), each test input at inference time is mapped to a belief function Shafer76 on the collection of classes $\mathcal{C}$. This belief function is encoded by a mass value $m(A)\in[0,1]$ assigned to a finite budget $\mathcal{O}=\{A_{k}\subset\mathcal{C}\}$ of sets of classes. The most relevant such sets are identified from training data by fitting Gaussian Mixture Models (GMMs) to the labelled data and computing the overlap among the resulting clusters spruyt2014draw . Such a predicted belief function is mathematically equivalent to a convex set of probability vectors (a credal set levi80book ; cuzzolin2008credal ) on the class list $\mathcal{C}$. A pointwise prediction can then be obtained by computing the centre of mass of this credal set, termed pignistic probability smets1994transferable . In RS-NN, uncertainty can be expressed either via the entropy of the pignistic prediction, or via the width of the credal set prediction antonucci2010credal . As we show in this paper, the entropy of a pignistic prediction is more diverse than that of a Bayesian prediction across in-distribution and out-of-distribution data (see Tab. 1), and the width of the credal prediction, as it encodes epistemic uncertainty about the prediction itself, is empirically much less correlated with the confidence score than entropy (see §D.5.1).
@@
Using a random-set representation prevents the need to discard information, a challenge observed in Bayesian Model Averaging hinne2020conceptual ; graefe2015limitations . While Bayesian inference requires defining prior distributions for model parameters even in the absence of relevant information, in belief theory, priors are not required for the inference process, thus avoiding the selection bias risks that can seriously condition Bayesian reasoning freedman1999wald . Based on our extensive experiments on multiple datasets, including large-scale ones, RS-NNs not only demonstrate superior accuracy compared to state-of-the-art Bayesian models (Sec. 4.2), but also arguably better encode the epistemic uncertainty associated with the predictions (Sec. 4.4) and better distinguish between in-distribution and out-of-distribution data (Sec. 4.3). Furthermore, RS-NNs effectively circumvent the tendency of standard neural networks to generate overconfident incorrect predictions, as illustrated in Fig. 2 in an experiment on Fast Gradient Sign Method (FGSM) goodfellow2014explaining adversarial attacks on CIFAR-10. CNN misclassifies with high confidence scores of 92.42% and 97.78%, while RS-NN shows lower confidence at 35.92% and 39.58%.
@@
Contributions. Firstly, we propose a novel Random-Set Neural Network (RS-NN) based on the principle that a deep neural network predicting belief values for sets of classes, rather than individual classes, has the potential to be a more faithful representation of the epistemic uncertainty induced by the limited quantity and quality of the training data. As they assign probability values to sets of outcomes directly, random sets Molchanov05 ; Molchanov_2017 naturally model the fact that observations almost invariably come in the form of sets, and can accommodate ambiguity, incomplete data, and non-probabilistic uncertainties. Secondly, we develop a novel budgeting method for efficiently selecting from the training data a fixed budget of relevant sets of classes for the task at hand, by fitting Gaussian Mixture Models to classes and computing clusters of classes, thus avoiding the exponential complexity of vanilla random-set implementations and ensuring scalability. A random-set representation serves as a foundational framework: when integrated with a budgeting of sets, this method acts as a versatile wrapper applicable to any model architecture. Thirdly, we illustrate the advantages of two new methods for assessing the uncertainty associated with a random-set prediction: the Shannon entropy of the pignistic prediction (the centre of mass of the corresponding credal set) and the width of the credal prediction itself. Finally, we present a large body of experimental results demonstrating how RS-NN outperforms both state-of-the-art Bayesian (BNN-R, LB-BNN) and Ensemble (ENN) methods in terms of: (i) performance (test accuracy, inference time) (Sec. 4.2); (ii) results on various out-of-distribution (OoD) benchmarks (Sec. 4.3), including CIFAR-10 vs. SVHN/Intel-Image, MNIST vs. FMNIST/KMNIST, and ImageNet vs. ImageNet-O; (iii) ability to provide reliable measures of uncertainty quantification (Sec. 4.4) in the form of pignistic entropy and credal set width, verified on OoD benchmarks; (iv) scalability to large-scale architectures (WideResNet-28-10, Inception V3, EfficientNet B2, ViT-Base) and datasets (e.g. ImageNet) (Sec. 4.5). Additionally, we provide statistical guarantees on RS-NN using non-conformity scores (§A), we show how RS-NN is robust to adversarial attacks (§D.2) and noisy data (§D.1), and circumvents the overconfidence problem in CNNs (§D.3). A qualitative assessment of entropy vs credal set width is given in §D.5.1.
@@
Paper outline. We first recall the notion of random sets, the mathematics of belief functions, and the associated credal and pignistic predictions (Sec. 2). We then explain the architecture, loss function and uncertainty representation of RS-NN (Sec. 3). In Sec. 4, we provide a large body of evidence supporting our approach. Sec. 4.6 details the limitations our method, including the budgeting time and fixed number of focal sets. Finally, Sec. 5 concludes and outlines future work. Appendix §A includes a discussion on statistical guarantees, §B, §C and §D contain additional evidence.
@@
Related Work. The machine learning community has recognised the challenge of estimating uncertainty in model predictions, leading to the development of various methods such as Bayesian approximations gal2016dropout , evidential Dirichlet distributions sensoy and conformal prediction shafer2008tutorial ; 10.5555/1712759.1712773 ; balasubramanian2014conformal ; crossconformal ; DBLP:journals/corr/abs-2107-07511 . Some methods rely on prior knowledge https://doi.org/10.48550/arxiv.2105.06868 , whereas others require setting a desired threshold on the metrics DBLP:journals/corr/abs-2107-07511 . Some baron1987second ; hullermeier2021aleatoric argued that classical probability theory is, in fact, not equipped to model “second-level" uncertainty on the probabilities themselves. This has led to the formulation of numerous uncertainty calculi cuzzolin2020geometry , including possibility theory Dubois90 , probability intervals halpern03book , credal sets levi80book , random sets Nguyen78 and imprecise probability walley91book .
@@
Bayesian approaches, pioneered by Buntine Buntine1991BayesianB and others 10.1162/neco.1992.4.3.448 ; neal2012bayesian , have recently become dominant in uncertainty estimation. Notable techniques include R-BNN (Reparameterisation) https://doi.org/10.48550/arxiv.1506.02557 , variational inference with reparameterisation trick and Laplace Bridge Bayesian approximation (LB-BNN) hobbhahn2022fast which uses the Laplace Bridge to map efficiently between Gaussian and Dirichlet distributions and enhance computational efficiency. Various approximations of full Bayesian inference exist, such as Markov chain Monte Carlo (MCMC), variational inference (VI), and Dropout Variational Inference https://doi.org/10.48550/arxiv.1506.02158 . Recent work addresses challenges in computational cost, model priors, and tractable function-space variational inference pmlr-v180-hobbhahn22a ; NEURIPS2021_a7c95857 ; tran2020all ; rudner2022tractable . Despite their advantages, Bayesian models face challenges when the model prior is misspecified. Our approach mitigates these challenges by eliminating the need for sampling during inference and prior selection, reducing computational complexity compared to Bayesian inference, as demonstrated in our experiments (see Tab. 2). Ensemble-based approaches, such as Deep Ensembles lakshminarayanan2017simple and Epistemic Neural Networks (ENN) osband2024epistemic , efficiently estimate uncertainty by leveraging multiple models. However, the computational cost of training ensembles, especially for large models, is often impractical (see Appendix §B).
@@ 2 Random sets and belief functions
Figure 3: The random set (set-valued random variable) associated with a cloaked die in which faces 1 and 2 are not visible.
Figure 4: A belief function is equivalent to a credal set with boundaries determined by lower bounds (4) on probability values.
Mathematically, this is a random set Matheron75 ; Kendall74foundations ; Nguyen78 ; Molchanov05 , i.e., a set-valued random variable, modelling random experiments in which observations come in the form of sets.
Belief functions. Random sets have been proposed by Dempster dempster2008upper and later Shafer Shafer76 as a mathematical model for subjective belief, alternative to Bayesian probability. For this reason, on finite domains (classification) they assume the name of belief functions. While classical discrete mass functions assign normalised, non-negative values to elements $\theta\in\Theta$ of their sample space, a belief function independently assigns normalised, non-negative mass values to subsets of the sample space:
$m(A)\geq 0,\forall A\subset\Theta,\sum_{A}m(A)=1.$ (1)
The belief function associated with a mass function $m$ then measures the total mass of the subsets of each focal set $A$. Mass functions can be recovered from belief functions via Moebius inversion Shafer76 :
$Bel(A)=\sum_{B\subseteq A}m(B),\>\>\>\>\>m(A)=\sum_{B\subseteq A}{(-1)^{|A\setminus B|}{Bel}(B)}.$ (2)
@@
Pignistic probability. Given a belief function $Bel$, its pignistic probability is the precise probability distribution obtained by re-distributing the mass of its focal sets $A$ to its constituent elements, $c\in A$:
$BetP(c)=\sum_{A\ni c}\frac{m(A)}{|A|}.$ (3)
Smets smets2005decision originally proposed to use the pignistic probability for decision making using belief functions, by applying expected utility to it. Notably, the pignistic probability is geometrically the centre of mass of the credal set associated with a belief function (cuzzolin2018visions , Fig. 4).
Credal prediction. As anticipated, RS-NN is designed to predict a belief function (a finite random set) on the set of classes. A belief function, in turn, is associated with a convex set of probability distributions (a credal set levi80book ; zaffalon-treebased ; cuzzolin2010credal ; antonucci10-credal ; antonucci2010credal ; cuzzolin2008credal ) on the same domain. This is the set:
$Cre=\left\{P:\mathcal{C}\rightarrow[0,1]|Bel(A)\leq P(A)\right\},$ (4)
where $P$ is any probability distribution on the collection of classes $\mathcal{C}$. As they have belief values as lower bounds, credal sets induced by belief functions have a peculiar shape (see Fig. 4, right, for a case with 3 classes). Their vertices are induced by the permutations of the elements of the sample space (for us, the set of classes) Chateauneuf89 ; cuzzolin2008credal . Given one such permutation (e.g., $(c_{2},c_{4},c_{1},c_{3})$ for a set of 4 classes) the probability vector representing the corresponding vertex of the credal prediction assigns to each specific class the mass of all the focal sets containing that class, but not containing any class preceding it in the permutation order wallner2005 .
The size of the resulting credal prediction measures the extent of the related epistemic uncertainty (see Sec. 3.3), arising from lack of evidence. A credal prediction, by encompassing multiple potential distributions, reflects the model’s acknowledgment of this uncertainty. A wider credal set indicates higher uncertainty, as the model refrains from committing to a specific probability distribution due to limited or conflicting evidence. In contrast, a narrower credal set implies lower uncertainty, signifying a more confident prediction based on substantial, consistent evidence.

### v3

@@ Abstract
Machine learning is increasingly deployed in safety-critical domains where erroneous predictions may lead to potentially catastrophic consequences, highlighting the need for learning systems to be aware of how confident they are in their own predictions: in other words, ‘to know when they do not know’. In this paper, we propose a novel Random-Set Neural Network (RS-NN) approach to classification which predicts belief functions (rather than classical probability vectors) over the class list using the mathematics of random sets, i.e., distributions over the collection of sets of classes. RS-NN encodes the ‘epistemic’ uncertainty induced by training sets that are insufficiently representative or limited in size via the size of the convex set of probability vectors associated with a predicted belief function. Our approach outperforms state-of-the-art Bayesian and Ensemble methods in terms of accuracy, uncertainty estimation and out-of-distribution (OoD) detection on multiple benchmarks (CIFAR-10 vs SVHN/Intel-Image, MNIST vs FMNIST/KMNIST, ImageNet vs ImageNet-O). RS-NN also scales up effectively to large-scale architectures (e.g. WideResNet-28-10, VGG16, Inception V3, EfficientNetB2 and ViT-Base-16), exhibits remarkable robustness to adversarial attacks and can provide statistical guarantees in a conformal learning setting.
@@ 1 Introduction
Machine learning models often struggle to provide reliable predictions when confronted with unfamiliar data (Guo et al., 2017a; Ovadia et al., 2019; Minderer et al., 2021), may it be noisy samples (Papernot et al., 2016) deliberately designed to deceive models, or out-of-distribution data (OoD) beyond the model’s training distribution. An ideal learning system, in opposition, should be aware of how confident it is in its own predictions, acknowledge the limits of its knowledge and gauge these limitations to make informed decisions by modelling the epistemic uncertainty associated with it, in essence, ‘to know when it does not know’. Epistemic uncertainty pertains to uncertainty associated with our ignorance about the underlying data generation process (Kendall & Gal, 2017; Hüllermeier & Waegeman, 2021; Manchingal & Cuzzolin, 2022). Within machine learning, a major source of epistemic uncertainty arises from the limited representativeness of the available training data, constrained in both quantity and quality.
@@
In this paper, we propose a new approach to classification which models epistemic uncertainty using a random set approach (Molchanov, 2005; 2017). As they assign probability values to sets of outcomes directly, random sets can naturally model the fact that observations often come in the form of sets (in particular, when missing data occurs), and accommodate ambiguity, incomplete data, and non-probabilistic uncertainties. As classification involves only a finite list of classes, we model uncertainty using belief functions (Shafer, 1976), the finite incarnation of random sets (Cuzzolin, 2018), whose theory (Shafer, 1976) is, in fact, a generalisation of Bayesian inference (Smets, 1986). Classical (discrete) probabilities can be seen as special belief functions, and Bayes’ rule as a special case of the Dempster’s rule of combination (Dempster, 2008) originally proposed for aggregating belief functions. For readers less familiar with the topic, in §A.1 we recall the distinction between classical probabilities and belief functions and the way they handle uncertainty in more detail. Fig. 2 contrasts the inference processes of our Random-Set Neural Network (RS-NN) and of a Bayesian Neural Network (Jospin et al., 2022). In hierarchical Bayesian inference (top), a posterior distribution over the network’s weights is learned from a training set. At prediction time, a predictive distribution is generated in the target space (left) by sampling from this weight posterior (middle), which amounts to a second-order probability distribution there (Hüllermeier & Waegeman, 2021). A single (mean) prediction is then typically derived by Bayesian Model Averaging (BMA) (Hoeting et al., 1999) (right), while uncertainty is measured by the entropy of the mean prediction and the variance of the predictive distribution.
Figure 1: Inference in a Bayesian Neural Network (top) as opposed to a Random-Set Neural Network (bottom), with corresponding measures of uncertainty and their sources. The triangle represents the set of probability vectors (probability simplex) one can define on the target space (e.g., a set of 3 classes).
Figure 2: Confidence scores of RS-NN and CNN for FGSM adversarial attack for different perturbations ($\epsilon$ = 0.05, 0.01) on MNIST dataset.
@@
In a Random-Set Neural Network (Fig. 2, bottom), each test input at inference time is mapped to a belief function (Shafer, 1976) on the collection of classes $\mathcal{C}$. This belief function is encoded by a mass value $m(A)\in[0,1]$ assigned to a finite budget $\mathcal{O}=\{A_{k}\subset\mathcal{C}\}$ of sets of classes (left). The most relevant such sets are identified from training data by fitting a Gaussian Mixture Model (GMM) to the labelled data and computing the overlap among the resulting clusters (Spruyt, 2013). Such a predicted belief function is mathematically equivalent to a convex set of probability vectors (credal set (Levi, 1980; Cuzzolin, 2008)) on the class list $\mathcal{C}$ (middle). A pointwise prediction (right) can then be obtained by computing the centre of mass of this credal set, termed pignistic probability (Smets & Kennes, 1994). In RS-NN, uncertainty can be expressed using either the entropy of the pignistic prediction (analogously to the Bayesian case), or the width of the credal set prediction (Antonucci & Cuzzolin, 2010). We find empirically that the pignistic entropy better separates in-distribution (iD) and out-of-distribution (OoD) samples than Bayesian entropy does (see Tab. 2); the width of the credal prediction, as it encodes the epistemic uncertainty about the prediction itself, is empirically much less correlated with the confidence score than entropy (see §E.5.3).
@@
Using a random-set representation prevents the need to discard information, a challenge observed in Bayesian Model Averaging (Hinne et al., 2020; Graefe et al., 2015). While Bayesian inference requires defining prior distributions for model parameters even in the absence of relevant information, in belief theory priors are not required for the inference process, thus avoiding the selection bias risks that can seriously condition Bayesian reasoning (Freedman, 1999). Based on our extensive experiments on multiple datasets, including large-scale ones, RS-NNs not only demonstrate superior accuracy compared to state-of-the-art Bayesian and Ensemble models (Sec. 4.2), but also arguably better encode the epistemic uncertainty associated with the predictions (Sec. 4.4) and better distinguish in-distribution and out-of-distribution data (Sec. 4.3). Furthermore, RS-NNs effectively circumvent the tendency of standard networks to generate overconfident incorrect predictions, as illustrated in Fig. 2 in an experiment on Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014) adversarial attacks on MNIST (LeCun & Cortes, 2005). CNN misclassifies with high confidence scores of 99.01% and 96.74%, while RS-NN shows lower confidence at 32.59% and 48.25%.
@@
Contributions. Firstly, we propose a novel Random-Set Neural Network (RS-NN) approach based on the principle that a deep neural network predicting belief values for sets of classes, rather than individual classes, has the potential to be a more faithful representation of the epistemic uncertainty induced by the limited quantity and quality of the training data. RS-NN is a ‘wrapper’ technique that can be applied on top of any existing baseline network model, by just changing the output layers and loss function. Statistical guarantees can also easily be provided for RS-NN predictions by applying conformal prediction on the pignistic probabilities (see §A.4). Secondly, we outline a budgeting method for efficiently selecting a limited budget of relevant sets of classes for the task at hand given the available training data, by fitting Gaussian Mixture Models to the labelled training points and computing their clusters. This overcomes the exponential complexity of vanilla random-set implementations, ensures the scalability of the approach to large datasets, and helps the network learn by limiting the available degrees of freedom (§E.7). Thirdly, we introduce two new methods for assessing the uncertainty associated with a random-set prediction: the Shannon entropy of the pignistic prediction and the width of the credal prediction itself, which prove to be more robust uncertainty measures. For instance, pignistic entropy provides a clearer distinction between in-distribution (iD) and out-of-distribution (OoD) entropy (Fig. 18, §E.5.1), while credal set width excels in separating iD and OoD samples, particularly for challenging datasets like ImageNet vs. ImageNet-O (Tab. 7). Finally, we present a large body of experimental results (based on a fair comparison principle in which all competing models are trained from scratch) which demonstrate how RS-NN outperforms both state-of-the-art Bayesian (LB-BNN (Hobbhahn et al., 2022), FSVI (Rudner et al., 2022)) and Ensemble (DE (Lakshminarayanan et al., 2017), ENN (Osband et al., 2024)) methods in terms of: (i) performance (test accuracy, inference time) (Sec. 4.2); (ii) results on various out-of-distribution (OoD) benchmarks (Sec. 4.3), including CIFAR-10 vs. SVHN/Intel-Image, MNIST vs. FMNIST/KMNIST, and ImageNet vs. ImageNet-O; (iii) ability to provide reliable measures of uncertainty quantification (Sec. 4.4) in the form of pignistic entropy and credal set width, verified on OoD benchmarks; (iv) scalability to large-scale architectures (WideResNet-28-10, Inception V3, EfficientNet B2, ViT-Base-16) and datasets (e.g. ImageNet) (Sec. 4.5). Additionally, we show how RS-NN is robust to adversarial attacks (§E.4) and noisy data (§E.3), and circumvents the overconfidence problem in CNNs (§E.2). A qualitative assessment of entropy vs credal set width is given in §E.5.3 and in-distribution vs out-of-distribution entropy scores are shown in Fig. 18.
@@
Paper outline. We first recall the notions of random sets, belief functions, credal and pignistic predictions (Sec. 2). We explain the RS-NN approach, loss function and uncertainty representation in Sec. 3. Sec. 4 provides a large body of empirical evidence supporting our approach and discusses its limitations. Sec. 5 concludes and outlines future work. Appendix §A discusses RS-NN learning process and statistical guarantees, §B reviews further related work, §C describes all algorithms in detail while §E contains a wealth of additional experimental results and all implementation details.
@@
Related Work. The machine learning community has recognised the challenge of estimating uncertainty in model predictions, leading to the development of several Bayesian approximations (Gal & Ghahramani, 2016; Charpentier et al., 2020), evidential Dirichlet models (Sensoy et al., 2018; Gao et al., 2024) and conformal prediction (Shafer & Vovk, 2008; Papadopoulos et al., 2008; Balasubramanian et al., 2014; Vovk, 2012; Angelopoulos & Bates, 2021). Some methods rely on prior knowledge (Fortuin, 2022), whereas others require setting a desired threshold on predictions (Angelopoulos & Bates, 2021). Some (Baron, 1987; Hüllermeier & Waegeman, 2021) have argued that classical probability is not equipped to model ‘second-level’ uncertainty on the probabilities themselves. This has led to the formulation of numerous uncertainty calculi (Cuzzolin, 2020), including possibility theory (Dubois & Prade, 1990), probability intervals (Halpern, 2017), credal sets (Levi, 1980), random sets (Nguyen, 1978) or imprecise probability (Walley, 1991).
@@
Bayesian approaches, pioneered by Buntine & Weigend (1991) and others (MacKay, 1992; Neal, 2012), are dominant in uncertainty estimation. Notable techniques include R-BNN (Reparameterisation) (Kingma et al., 2015), variational inference with reparameterisation and Laplace Bridge Bayesian approximation (LB-BNN) (Hobbhahn et al., 2022), which uses the Laplace Bridge to map between Gaussian and Dirichlet distributions. Various approximations of full Bayesian inference exist, such as Markov chain Monte Carlo (MCMC), function-space BNNs (Sun et al., 2019) such as function-space variational inference (FSVI) (Rudner et al., 2022), and Dropout Variational Inference (Gal & Ghahramani, 2015). In our experiments, we do not consider older Bayesian models such as R-BNN, MCMC and Dropout VI, since they have been superseded by more performing approaches and are computationally expensive to train on larger datasets and architectures. Recent work addresses challenges in computational cost (Hobbhahn et al., 2022; Daxberger et al., 2021) and model priors (Tran et al., 2020). Despite their advantages, Bayesian models face challenges when the model prior is misspecified. Ensemble-based approaches, such as Deep Ensembles (DE) (Lakshminarayanan et al., 2017) and Epistemic Neural Networks (ENN) (Osband et al., 2024), efficiently estimate uncertainty by leveraging multiple models. However, the computational cost of training ensembles, especially for large models, is often impractical. Our approach mitigates these challenges by eliminating the need for both inference-time sampling and prior selection, reducing computational complexity compared to Bayesian inference and lowering training time compared to Ensembles, as demonstrated in our experiments in Sec. 4. More related work is given in §B.
@@ 2 Random sets and belief functions
Figure 3: The random set associated with a cloaked die in which faces 1 and 2 are not visible.
Figure 4: A belief function is equivalent to a credal set with boundaries determined by lower bounds (Eq. 4) on probability values.
Belief functions. Random sets have been proposed by Dempster (2008) and Shafer (1976) as a mathematical model for subjective belief, alternative to Bayesian reasoning. Thus, on finite domains (e.g., for classification) they assume the name of belief functions. While classical discrete mass functions assign normalised, non-negative values to elements $\theta\in\Theta$ of their sample space, a belief function independently assigns normalised, non-negative mass values to its subsets $A\subset\Theta$:
$m(A)\geq 0,\>\>\>\>\>\sum_{A}m(A)=1,\>\>\forall A\in\mathbb{P}(\Theta)\vspace{-6pt}$ (1)
The belief function associated with a mass function $m$ measures the total mass of the subsets of each ‘focal set’ $A$. Mass functions can be recovered from belief functions via Moebius inversion (Shafer, 1976), which, in combinatorics, plays a role similar to that of the derivative:
$Bel(A)=\sum_{B\subseteq A}m(B),\>\>\>\>\>m(A)=\sum_{B\subseteq A}{(-1)^{|A\setminus B|}{Bel}(B)}.\vspace{-6pt}$ (2)
@@
Pignistic probability. Given a belief function $Bel$, its pignistic probability is the precise probability distribution obtained by re-distributing the mass of its focal sets $A$ to its constituent elements, $\theta\in A$:
$BetP(\theta)=\sum_{A\ni\theta}\frac{m(A)}{|A|}.$ (3)
Smets (2005) originally proposed to use the pignistic probability for decision making using belief functions, by applying expected utility to it. Notably, the pignistic probability is geometrically the centre of mass of the credal set (see Fig. 4) associated with a belief function (Cuzzolin, 2018).
Credal prediction. As anticipated, RS-NN is designed to predict a belief function (a finite random set) on the set of classes. A belief function, in turn, is associated with a convex set of probability distributions (a credal set (Levi, 1980; Zaffalon & Fagiuoli, 2003; Cuzzolin, 2010; Antonucci & Cuzzolin, 2010; Cuzzolin, 2008)) on the same domain. This is the set:
$Cre=\left\{P:\Theta\rightarrow[0,1]|Bel(A)\leq P(A)\right\},$ (4)
of probability distributions $P$ on $\Theta$ which dominate the belief function on each focal set $A$. The size of the resulting credal prediction measures the extent of the related epistemic uncertainty (see Sec. 3.3), arising from lack of evidence (see §E.5.3). The use of credal set size as a measure of epistemic uncertainty is well-supported in literature (Hüllermeier & Waegeman, 2021; Bronevich & Klir, 2008), as it aligns with established concepts of uncertainty such as conflict and non-specificity (Yager, 2008; Kolmogorov, 1965). A credal prediction, by encompassing multiple potential distributions, reflects the model’s acknowledgment of this uncertainty. A wider credal set indicates higher uncertainty, as the model refrains from committing to a specific probability distribution due to limited or conflicting evidence. In contrast, a narrower credal set implies lower uncertainty, signifying a more confident prediction based on substantial, consistent evidence.